OpenStudy (anonymous):

If y is 6 when x is 3 and y varies directly with x, find y when x is 8. a) 4 b) 11 c) 16

OpenStudy (godlovesme):

y directly varies with x means \[y/x\]

OpenStudy (godlovesme):

so substitute the numbers they gave u for y and x 6/3 and y/8

OpenStudy (godlovesme):

@makayla.boze make sense?

OpenStudy (anonymous):

yes, I get it now @Godlovesme

OpenStudy (godlovesme):

great :) can u solve for y ?

OpenStudy (godlovesme):

6/3= x/8 first do cross multiplication

OpenStudy (pawanyadav):

Let y=kx K is proportionality constant. As. 6=k×3 k=6/3=2 Now Y=k×8 As k=2 Y=2×8=16

OpenStudy (anonymous):

ok, that's what I thought @Godlovesme

OpenStudy (pawanyadav):

Did that help you

OpenStudy (godlovesme):

@Pawanyadav nice job but please let the users do some of the work too :)

OpenStudy (anonymous):

thanks @Godlovesme

OpenStudy (godlovesme):

welcome (: